OpenStudy (anonymous):

6/20 as a percent is.......

OpenStudy (anonymous):

30%

OpenStudy (anonymous):

6/20=X/100 X=30

OpenStudy (asnaseer):

to convert any fraction or decimal to a percentage just multiply it by 100. e.g. 3/4 is:\[\frac{3}{4}*100=75\%\] e.g. 0.45 is:\[0.45*100=45\%\]